# The Rise and Fall of the Initial Era

### v1

@@ I Introduction
In the contemporary discourse on technology, dominated by references to digital and electronic innovations, the notion of the research article as a form of technology may appear incongruous. However, an exploration of the research article through the lens of technology is critical for a comprehensive understanding of its role, interactions with other technological forms, and its consequent impact on society. The research article, viewed technologically, is a significant construct, with a long-standing history of shaping social norms and establishing institutions that extend their influence across the research community, irrespective of disciplinary boundaries, geographical locations, or historical periods Hall (2001). This technology has achieved ubiquity on three distinct levels: spatial, with researchers globally engaging with and understanding research articles under shared assumptions; temporal, allowing for the contextual comprehension of older articles through a slow evolution of the format; and disciplinary, with a cross-disciplinary recognition of the rigorous scrutiny and scientific methodology underpinning the work. These characteristics are critical for research to function, as an incremental activity that builds on prior results and knowledge.
@@
[…]
The study of the representation of author names provides us with an insight into the structure of research culture and the sociology of research. This has been well-studied with the implications of gender bias in various aspects of the scholarly record being examined in detail Macaluso et al. (2016); Ni et al. (2021); Sugimoto et al. (2019); Kozlowski et al. (2022a, b). It is not the focus of this article to rehash these arguments but rather to add to the data supporting these arguments. In a period where there are high proportions of researchers using full-form names, we are able potentially to determine more about the gender and background of the participants in the research ecosystem and thus learn about the geographical and discipline-based diversity of the participants in the research ecosystem. When we enter periods where initial-form is used, data become less rich and we are able to determine less.
This study leverages the slow evolution of the research paper to conduct a form of “bibliometric archaeology” Hook and Porter (2021); Bornmann et al. (2021), examining the interplay between technological and societal changes as encapsulated in the history of the research paper’s format. By focusing on seemingly minor details, such as the presentation of author names, from full given names (“full form”) to initials only (“initial form”), this paper seeks to uncover broader narratives, correlating shifts in norms with technological, geographic, disciplinary, and societal factors.
@@ I.1 A brief history of metadata norms
Figure 1: A page from the Philosophical Transactions of the Royal Society from December 1669 demonstrating at once the similarity to a modern research article and the significant differences in the level and detail of metadata present in the article noa (1997).
@@
The letter is dated but not signed and the author is not explicitly named noa (1997).
@@
The reasons for omitting author identities can also stem from more problematic motives, such as the desire to conceal an author’s gender. Now famously, the Brontë sisters initially published under male pseudonyms as writing was not an appropriate activity for women in that period. More recently J.K.Rowling opted to use her initials to conceal her gender for commercial reasons, anticipating potential bias in her readership, before she became too well known and successful to be able to conceal her identity Whited (2004).
Anonymity or pseudonymity can be a choice not only to protect from persecution, it can also serve to protect authors from repercussions related to the voicing of controversial opinions or work that is not in step with the political regime where it is being carried out. The revolutionary ideas shared on the future of currency are an example in the seminal white paper on Bitcoin by Satoshi Nakamoto, where the identity, gender, nationality, or even the individuality or plurality of the author(s) remains unknown Nakamoto (2008).
@@ I.2 The evolution of the research article as technology
The role of technology in facilitating or restricting the disclosure of author information has pivotal import. For example, while current practices do not mandate the encoding of gender data in author identities, this omission has been significant in fields where research outcomes vary with the researcher’s gender Sorge et al. (2014); Katsnelson (2014); Reardon (2017). The decision not to maintain detailed records of authors’ gender or ethnicity is influenced by contemporary societal norms, yet future scholars might view this lack of information as a critical oversight, akin to how the absence of affiliation details in early scholarly works is perceived today.
@@
The interplay between technological and societal changes, each influencing the other in complex ways, is also evident in the shifting patterns of co-authorship in scholarly papers. Historical analysis (see Fig. 2) reveals a shift in the modal average number of authors from one to gradually increasing numbers of participants over time, reflecting the changing sociology of research. Analysis of the locations of researchers shows the fundamental nature of research community Adams (2013).
The slow and steady evolution shown in Fig. 2 in collaboration may also be considered to be a societal parallel for the slow evolution in the nature of the technologies that underpin scholarly communication. The increase in authorship of a paper from a single author (modal average) from the genesis of scholarly communication until around 1960 followed by a shift, relatively rapidly, to five co-authors per paper as a modal average in the present day, gives us a sense of when technological and social changes started to facilitate and change norms.
@@
[…]
Illustrative of this trend, the growth in the number of research papers featuring extensive co-authorship is striking (see Fig. 3). Papers with more than 100 authors, indicative of large collaborative projects, began to appear with greater frequency, reflecting the complex, interdisciplinary nature of modern scientific inquiries (yellow region) starts in earnest in the late 1970s [^1] The first paper in Dimensions with co-authorship of more than 100 authors is in Chemistry from 1928 Emich et al. (1928).. This trend is further accentuated in research requiring substantial resources, such as particle physics experiments, where the collective expertise and effort of hundreds of scientists are essential to the project’s success. Papers with over 500 co-authors (blue) and, notably, more than 1000 co-authors (red) have become increasingly common, underscoring the scale of collaboration in certain areas of research.
@@
On a practical level, the significance of author names extends beyond mere attribution. In the context of evaluations for promotions, tenure, grant assessments, and other academic recognition, the identity of authors, corroborated by associated identifiers, plays a crucial role. The trust placed in a particular paper, and by extension, in communities of researchers, is increasingly dependent on the clarity and reliability of authorship information. Furthermore, this information is pivotal in selecting referees, assessing potential conflicts of interest, and facilitating a myriad of other critical academic processes. The evolution of scholarly communication, therefore, not only reflects changes in the academic landscape but also emphasises the growing importance of clear and reliable attribution in maintaining the integrity and trustworthiness of scholarly work.
@@ II Methodology
In the study presented here, we explore the change in the representation of author names in scholarly research articles. As established in the Introduction, we define the terms initial form to refer to an author name where only the initials are supplied and no given name is supplied, and full form to refer an author name that includes at least one fully stated given name. We use Dimensions from Digital Science as the data source for our exploration. At the time of writing Dimensions contains over 144 million publications, of which approximately 112 million are journal articles. Much of the analysis shown in this paper is based on queries that can be run directly on the Dimensions database in the Google BigQuery environment Hook and Porter (2021).
[…]
@@
In terms of both the quality and volume of the data, the transition from the anecdotal to the statistical epoch is not sharply defined but happens gradually. However, our choice of 1945 is not completely arbitrary. The current authors have previously noted that 1945 was the year in which the centre of mass of global research stopped its march toward North America and began to move back toward Europe Hook and Porter (2021). It was also the year in which Vannevar Bush published the Endless Frontier Bush (1945) and, as such, is a good point in time from which to date the modern approach to research. Indeed, both Figs 2 and 4 show a marked change in behaviour between 1940 and 1950.
@@ 20    UNNEST(authors) author
It would be a simple matter to test the first name to see if it is of length one or two characters (either an initial or an initial and a “.”). However, this approach would miss people with multiple initials or with a two character first name such as “Bo”. Hence, we use a slightly more complex approach were we examine the first three segmented tokens in the first name field in Dimensions. We test for the length of the field, the position of the character “.” . This allows us to correctly classify edge cases such as authors that are known principally, and hence who publish by, their second name such as J. Robert Oppenheimer (see, for example Oppenheimer (1963)).
A further complexity is that, in many cases, names are preserved in both their original and transliterated forms in Dimensions. In these cases a whole name is a single character, for example in Kanji or Katakana (see, for example Yan et al. (2005)) and hence a whole name can be mistaken for an initial. Fortunately, this problem is a small one in the context of the current work since, as of the date of writing, only 250,991 of 8,456,140 publications in Chinese (or around 3%) have the potential to contain Kanji characters, that might have have been transliterated into English. There is a potentially bigger issue for Japan in that 1,152,602 of 4,460,000 Japanese language publications (or around 25%) have the potential to contain Kanji or Katakana characters without transliteration. However, deeper examination suggests that just 176,177 of 2,741,338 authors (or around 6.5%) with a Japanese address have a single character first name - which may legitimately be a transliterated first initial or a Japanese character that could be misclassified as an initial. While this number is not insignificant, we only wish to pick out high-level trends in the data in this article and hence this is negligible. We have not commented on how the percentage of Chinese language and Japanese language publications requiring translation has changed with time, however, this point is currently moot as global research has been dominated by countries with roman alphabets in the past and, again, for an analysis that requires only a high-level trend, it is reasonable to assume that the variance introduced by this effect does not change the overall appearance of the line.
@@
In exploring the issue of gender we used genderize.io under licence. The genderize.io tool provides a dataset with probabilities that a given name is statistically associated with either men or women, or “undetermined” on both a nation-neutral and national basis. This is only a statistical analysis and hence cannot attempt to access the subtle and complex issues of gender identity, but rather it focuses on how gender is represented in the historical scholarly record. We have employed a methodology where we matched name variations in author names in Dimensions to the results obtained for that name variation in genderize.io. The results were stored in a private table in Google BigQuery. We then matched author names, where given, to data in this table. We used the current university of the academic to provide nationality information and, where data was either unavailable or indeterminate, we defaulted to using data without a national context. Obviously, this approach is open to errors in that, especially in more recent years, researchers are much more likely to travel during their careers and hence their current university may not be culturally relevant for such an analysis. Many names are also used much more internationally now and hence those that are traditionally nationally associated may not be used in context. It is also the case that some researchers will originate from countries where a Western name is adopted for business purposes. In this final case, the names are often more central to common usage as they are designed to mimic Western-style naming conventions and may even improve the data. The point of this paper is not, however, to do a detailed gender analysis and we believe that our approach is sufficiently robust to be serviceable in the current context.
@@ III.1 Paper-based analysis
As indicated in the introduction, the historical use of names in the scholarly record has been uneven and may, to a certain extent, reflect societal norms. While, in the early years of Dimensions’ coverage, the volumes of data are small, Fig. 5 shows that until around 1800 it appeared to be normal to use one’s full name in scholarly correspondence (despite the fascinating cases discussed in the introduction). It is also important to acknowledge the lens through which we are examining the past is imperfect. The work of Lefanu Lefanu (1937) makes it evident that there was a lively and thriving research discussion prior to 1800, however, records have either not been preserved or not been transferred into digital realms, meaning that they are invisible to our study.
@@
  • We define the “Developmental Era” as the period from 1665 to 1950, that appears to be split into two sections, separated by a sharp transition that occurs in 1798. As we will explain, the sharpness of this distinction is a result of the data in Dimensions (and of the systems on which Dimensions is based).
    In the period until 1798 there were just a few regularly publishing journals with total global number of articles being published reaching 288 in 1798. Up until that year, the journals with the largest publication volumes were Philosophical Transactions of the Royal Society of London, Transactions of the Linnean Society of London and Transactions of the Royal Society of Edinburgh, each of which were publishing around 30 articles per year and all of which had adopted a standardised title page, typically printed the names of authors in full form. But, in 1798, The Philosophical Magazine: A Journal of Theoretical, Experimental and Applied Physics was established by Alexander Tilloch of the London Philosophical Society, publishing 142 articles; the following year, in 1799, an even more voluminous journal launched as The British and foreign medical review was launched by the Royal College of Physicians, publishing 327 articles. Hence, there was a sharp and significant increase in the number of articles in Dimensions. These are the articles in publications that scholarly societies and other participants in the scholarly ecosystem decided were sufficiently valuable (and where means were sufficient) to clean up metadata and make DOIs or PubMed identifiers available. In the two significant journals mentioned here, author name forms are mixed in both journals with some articles being editorial (without author names), but while full form names still appear regularly, initial form becomes much more acceptable and this becomes established as a dominant signal entering the 1800s. Lefanu provides a for more detailed treatment of the rise of medical journals Lefanu (1937).
@@
    It is worthy of note that throughout the whole of the Developmental Era there a low level of co-authorship and hence, we remain in Adams’ first age Adams (2013). As we see in Fig. 4 it is not until 1958 that the modal average of authors on a paper moves from single-authorship to two co-authors.
@@
  • The “Initial Era” began in the wake of the second world war, the dawn of Bush’s Endless Frontier Bush (1945) in 1945, and continues until 1980. It is the singular period in the history of the scholarly record in which the initial form dominated the full form of author names. The initial era is entirely contained in the statistical epoch and hence there are sufficiently many journals and papers that a statistical approach to analysis is valid.

### v2

@@ I Introduction
In the contemporary discourse on technology, dominated by references to digital and electronic innovations, the notion of the research article as a form of technology may appear incongruous. However, an exploration of the research article through the lens of technology is critical for a comprehensive understanding of its role, interactions with other technological forms, and its consequent impact on society. The research article, viewed technologically, is a significant construct, with a long-standing history of shaping social norms and establishing institutions that extend their influence across the research community, irrespective of disciplinary boundaries, geographical locations, or historical periods [1]. This technology has achieved ubiquity on three distinct levels: spatial, with researchers globally engaging with and understanding research articles under shared assumptions; temporal, allowing for the contextual comprehension of older articles through a slow evolution of the format; and disciplinary, with a cross-disciplinary recognition of the rigorous scrutiny and scientific methodology underpinning the work. These characteristics are critical for research to function, as an incremental activity that builds on prior results and knowledge.
@@
At its core, the research paper remains a rare example of a 17th-Century technology in current use. It is a technology that originally emerged from a very different time to serve a very different context. It was the Age of Empire, where research communications were characterised by brief correspondences, often containing hand-drawn diagrams or concise result tables, exchanged among a small, affluent elite. Originating in the club culture of coffee houses in cities like London and Paris, this technology now underpins a vast, global research ecosystem, supporting millions of contributors and a burgeoning diversity of cultures, geographies, and subjects [2, 3, 4].
The study of the representation of author names provides us with an insight into the structure of research culture and the sociology of research. This has been well-studied with the implications of gender bias in various aspects of the scholarly record being examined in detail [5, 6, 7, 8, 9]. It is not the focus of this article to rehash these arguments but rather to add to the data supporting these arguments. In a period where there are high proportions of researchers using full-form names, we are able potentially to determine more about the gender and background of the participants in the research ecosystem and thus learn about the geographical and discipline-based diversity of the participants in the research ecosystem. When we enter periods where initial-form is used, data become less rich and we are able to determine less.
This study leverages the slow evolution of the research paper to conduct a form of “bibliometric archaeology” [10] over the digital research record [11, 12], examining the interplay between technological and societal changes as encapsulated in the history of the research paper’s format. By focusing on seemingly minor details, such as the presentation of author names, from full given names (“full form”) to initials only (“initial form”), this paper seeks to uncover broader narratives, correlating shifts in norms with technological, geographic, disciplinary, and societal factors.
@@ I.1 A brief history of metadata norms
Figure 1: A page from the Philosophical Transactions of the Royal Society from December 1669 demonstrating at once the similarity to a modern research article and the significant differences in the level and detail of metadata present in the article [13].
@@
The letter is dated but not signed and the author is not explicitly named [13].
@@
The reasons for omitting author identities can also stem from more problematic motives, such as the desire to conceal an author’s gender. Now famously, the Brontë sisters initially published under male pseudonyms as writing was not an appropriate activity for women in that period. More recently J.K.Rowling opted to use her initials to conceal her gender for commercial reasons, anticipating potential bias in her readership, before she became too well known and successful to be able to conceal her identity [14].
Anonymity or pseudonymity can be a choice not only to protect from persecution, it can also serve to protect authors from repercussions related to the voicing of controversial opinions or work that is not in step with the political regime where it is being carried out. The revolutionary ideas shared on the future of currency are an example in the seminal white paper on Bitcoin by Satoshi Nakamoto, where the identity, gender, nationality, or even the individuality or plurality of the author(s) remains unknown [15].
@@ I.2 The evolution of the research article as technology
The role of technology in facilitating or restricting the disclosure of author information has pivotal import. For example, while current practices do not mandate the encoding of gender data in author identities, this omission has been significant in fields where research outcomes vary with the researcher’s gender [16, 17, 18]. The decision not to maintain detailed records of authors’ gender or ethnicity is influenced by contemporary societal norms, yet future scholars might view this lack of information as a critical oversight, akin to how the absence of affiliation details in early scholarly works is perceived today.
@@
The interplay between technological and societal changes, each influencing the other in complex ways, is also evident in the shifting patterns of co-authorship in scholarly papers. Historical analysis (see Fig. 2) reveals a shift in the modal average number of authors from one to gradually increasing numbers of participants over time, reflecting the changing sociology of research. Analysis of the locations of researchers shows the fundamental nature of research community [19].
The slow and steady evolution shown in Fig. 2 in collaboration may also be considered to be a societal parallel for the slow evolution in the nature of the technologies that underpin scholarly communication. The increase in authorship of a paper from a single author (modal average) from the genesis of scholarly communication until around 1960 followed by a shift, relatively rapidly, to five co-authors per paper as a modal average in the present day, gives us a sense of when technological and social changes started to facilitate and change norms. A more detailed analysis of evolution of co-authorship and its importance in addressing more complex problems is dealt with by Thelwall and Maflahi [20].
@@
The 20th century witnessed a transformative shift in research practices, fundamentally driven by technological advancements. This period marked the emergence of “Big Science”, a paradigm characterised by large-scale scientific endeavours that required extensive collaboration across disciplines and often, across national borders. The sociological landscape of certain research fields underwent significant changes due to this shift, influencing and reshaping established norms within the scientific community. The impact of Big Science is particularly evident in the evolving patterns of authorship, as collaborations expanded to include hundreds or even thousands of contributors.
Illustrative of this trend, the growth in the number of research papers featuring extensive co-authorship is striking (see Fig. 3). Papers with more than 100 authors, indicative of large collaborative projects, began to appear with greater frequency, reflecting the complex, interdisciplinary nature of modern scientific inquiries (yellow region) starts in earnest in the late 1970s  [^1] The first paper in Dimensions with co-authorship of more than 100 authors is in Chemistry from 1928 [66].. This trend is further accentuated in research requiring substantial resources, such as particle physics experiments, where the collective expertise and effort of hundreds of scientists are essential to the project’s success. Papers with over 500 co-authors (blue) and, notably, more than 1000 co-authors (red) have become increasingly common, underscoring the scale of collaboration in certain areas of research.
@@
On a practical level, the significance of author names extends beyond mere attribution. In the context of evaluation for promotions, tenure, grant assessments, and many other forms academic recognition, the identity of authors, corroborated by associated identifiers, plays a crucial role. Indeed, research is becoming increasingly centred around quantification [22]. The trust placed in a particular paper, and by extension, in communities of researchers, is increasingly dependent on the clarity and reliability of authorship information. Furthermore, this information is pivotal in selecting referees, assessing potential conflicts of interest, and facilitating a myriad of other critical academic processes. The evolution of scholarly communication, therefore, not only reflects changes in the academic landscape but also emphasises the growing importance of clear and reliable attribution in maintaining the integrity and trustworthiness of scholarly work.
@@ II Methodology
In the study presented here, we explore the change in the representation of author names in scholarly research articles. As established in the Introduction, we define the terms initial form to refer to an author name where only the initials are supplied and no given name is supplied, and full form to refer an author name that includes at least one fully stated given name. We use Dimensions from Digital Science as the data source for our exploration. At the time of writing Dimensions contains over 144 million publications, of which approximately 112 million are journal articles. Much of the analysis shown in this paper is based on queries that can be run directly on the Dimensions database in the Google BigQuery environment without the need for further coding in, for example, a Python environment and hence are accessible to a broad audience [11, 23].
Although we have included several graphs and examples regarding scholarly output before 1945 we recognise that publications volumes are smaller further back in time. Thus, we define two epochs in publication, one before 1945 which we refer to as the anecdotal epoch and the other after 1945 that we call the statistical epoch. It is noteworthy that Dimensions, our primary data source, is limited in the validity of its data for earlier times in the anecdotal epoch. The data source was not constructed with the current use case in mind and hence coverage of content in the 17th, 18th and even 19th centuries is reliant on publisher and community interest in awarding persistent identifiers to research outputs of these periods in order for them to be included in Dimensions. In addition, mappings to institutions that have disappeared or changed their name over the earlier part of the last 350 years may not be fully represented in geographical analyses - similarly, country references will be to modern countries not countries that have existed in the past - we take no account of moving boundaries, merely imposing the current world view back in time for simplicity of analysis.
@@
In terms of both the quality and volume of the data, the transition from the anecdotal to the statistical epoch is not sharply defined but happens gradually. However, our choice of 1945 is not completely arbitrary. The current authors have previously noted that 1945 was the year in which the centre of mass of global research stopped its march toward North America and began to move back toward Europe [11]. It was also the year in which Vannevar Bush published the Endless Frontier [24] and, as such, is a good point in time from which to date the modern approach to research. Indeed, both Figs 3 and 4 show a marked change in behaviour between 1940 and 1950.
@@ 20 UNNEST(authors) author
It would be a simple matter to test the first name to see if it is of length one or two characters (either an initial or an initial and a “.”). However, this approach would miss people with multiple initials or with a two character first name such as “Bo”. Hence, we use a slightly more complex approach were we examine the first three segmented tokens in the first name field in Dimensions. We test for the length of the field, the position of the character “.” . This allows us to correctly classify edge cases such as authors that are known principally, and hence who publish by, their second name such as J. Robert Oppenheimer (see, for example [25]).
A further complexity is that, in many cases, names are preserved in both their original and transliterated forms in Dimensions. In these cases a whole name is a single character, for example in Kanji or Katakana (see, for example [26]) and hence a whole name can be mistaken for an initial. Fortunately, this problem is a small one in the context of the current work since, as of the date of writing, only 250,991 of 8,456,140 publications in Chinese (or around 3%) have the potential to contain Kanji characters, that might have have been transliterated into English. There is a potentially bigger issue for Japan in that 1,152,602 of 4,460,000 Japanese language publications (or around 25%) have the potential to contain Kanji or Katakana characters without transliteration. However, deeper examination suggests that just 176,177 of 2,741,338 authors (or around 6.5%) with a Japanese address have a single character first name - which may legitimately be a transliterated first initial or a Japanese character that could be misclassified as an initial. While this number is not insignificant, we only wish to pick out high-level trends in the data in this article and hence this is negligible. We have not commented on how the percentage of Chinese language and Japanese language publications requiring translation has changed with time, however, this point is currently moot as global research has been dominated by countries with roman alphabets in the past and, again, for an analysis that requires only a high-level trend, it is reasonable to assume that the variance introduced by this effect does not change the overall appearance of the line.
@@
In exploring the issue of gender we used genderize.io under licence. The genderize.io tool provides a dataset with probabilities that a given name is statistically associated with either men or women, or “undetermined” on both a nation-neutral and national basis. The tool is well-used in the study of gender in academia [27, 28]. This is a statistical analysis and hence cannot attempt to access the subtle and complex issues of gender identity, but rather it focuses on how gender is represented in the historical scholarly record. We have employed a methodology where we matched name variations in author names in Dimensions to the results obtained for that name variation in genderize.io. The results were stored in a private table in Google BigQuery. We then matched author names, where given, to data in this table. We used the current university of the academic to provide nationality information and, where data was either unavailable or indeterminate, we defaulted to using data without a national context. Obviously, this approach is open to errors in that, especially in more recent years, researchers are much more likely to travel during their careers and hence their current university may not be culturally relevant for such an analysis. Many names are also used much more internationally now and hence those that are traditionally nationally associated may not be used in context. It is also the case that some researchers will originate from countries where a Western name is adopted for business purposes. In this final case, the names are often more central to common usage as they are designed to mimic Western-style naming conventions and may even improve the data. The point of this paper is not, however, to do a detailed gender analysis and we believe that our approach is sufficiently robust to be serviceable in the current context.
@@ III.1 Paper-based analysis
As indicated in the introduction, the historical use of names in the scholarly record has been uneven and may, to a certain extent, reflect societal norms. While, in the early years of Dimensions’ coverage, the volumes of data are small, Fig. 5 shows that until around 1800 it appeared to be normal to use one’s full name in scholarly correspondence (despite the fascinating cases discussed in the introduction). It is also important to acknowledge the lens through which we are examining the past is imperfect. The work of Lefanu [29] makes it evident that there was a lively and thriving research discussion prior to 1800, however, records have either not been preserved or not been transferred into digital realms, meaning that they are invisible to our study.
@@
  • We define the “Developmental Era” as the period from 1665 to 1950, that appears to be split into two sections, separated by a sharp transition that occurs in 1798. As we will explain, the sharpness of this distinction is a result of the data in Dimensions (and of the systems on which Dimensions is based–specifically, systems that are based around the attribution of DOIs to academic articles [30]). Thus, we must regard the data reported until 1798 as unrepresentative of the actual development of scholarly publication and only anecdotal in nature.
    In the period until 1798 the Dimensions data contains just a few regularly publishing journals with total global number of articles being published reaching 288 in 1798. Up until that year, the journals with the largest publication volumes were Philosophical Transactions of the Royal Society of London, Transactions of the Linnean Society of London and Transactions of the Royal Society of Edinburgh, each of which were publishing around 30 articles per year and all of which had adopted a standardised title page, typically printed the names of authors in full form. But, in 1798, The Philosophical Magazine: A Journal of Theoretical, Experimental and Applied Physics was established by Alexander Tilloch of the London Philosophical Society, publishing 142 articles; the following year, in 1799, an even more voluminous journal launched as The British and foreign medical review was launched by the Royal College of Physicians, publishing 327 articles. Hence, there was a sharp and significant increase in the number of articles in Dimensions. These are the articles in publications that scholarly societies and other participants in the scholarly ecosystem decided were sufficiently valuable (and where means were sufficient) to clean up metadata and make DOIs or PubMed identifiers available. In the two significant journals mentioned here, author name forms are mixed in both journals with some articles being editorial (without author names), but while full form names still appear regularly, initial form becomes much more acceptable and this becomes established as a dominant signal entering the 1800s. Lefanu provides a for more detailed treatment of the rise of medical journals [29].
@@
    It is worthy of note that throughout the whole of the Developmental Era there a low level of co-authorship and hence, we remain in Adams’ first age [19]. As we see in Fig. 2 it is not until 1958 that the modal average of authors on a paper moves from single-authorship to two co-authors.
@@
  • The “Initial Era” began in the wake of the second world war, the dawn of Bush’s Endless Frontier [24] in 1945, and continues until 1980. It is the singular period in the history of the scholarly record in which the initial form dominated the full form of author names. The initial era is entirely contained in the statistical epoch and hence there are sufficiently many journals and papers that a statistical approach to analysis is valid.